OpenStudy (yrelhan4):

The driver of a car moving on a straight road applies brakes to come to rest, with a constant retardation 'a'. Assuming that the time of motion is more than 2 s, the distance covered by the car in secondlast second of its motion is

OpenStudy (anonymous):

Distance Travelled in nth Second \[Dn = u + \frac{ a }{ 2 }[2n-1]\] Since the Question Talks about retardation \[Dn = u - \frac{ a }{ 2 }[2n-1]\]

OpenStudy (anonymous):

\[v = u -at\] \[u = at\] t - 1 = n t = n +1 u = a (n+1)

OpenStudy (anonymous):

Nw Substitute the Value of u in the 1st equation

OpenStudy (anonymous):

\[Dn = an + a - an + a/2 = 3a/2\]

OpenStudy (yrelhan4):

Thank you! (:

OpenStudy (anonymous):

Welcome...

OpenStudy (yrelhan4):

why is t-1=n?

OpenStudy (anonymous):

secondlast second of its motion

OpenStudy (anonymous):

let t be the total time

OpenStudy (anonymous):

secondlast = t-1

OpenStudy (yrelhan4):

oh.. haha.. ty again!